OpenStudy (calculusxy):

Which equation has a graph that contains points that slant downward, left to right, and crosses the y-axis at y =4? a. x + y = -4 b. y - x = -4 c. y - x = 4 d. x + y = 4 @iambatman @TheSmartOne

OpenStudy (calculusxy):

@iambatman I was going to post the question on the previous post, but it somehow did not work.

TheSmartOne (thesmartone):

So if the y-intercept is 4 then the line must have (0,4) as as solution...

TheSmartOne (thesmartone):

And A and D look the same...

OpenStudy (anonymous):

Do you know of the form y = mx+b? This is initially dealing with such a line. It's called the slope intercept form. Try seeing if you can put it in such a way and graph it out using some points.

OpenStudy (calculusxy):

What is y = mx + b ?

OpenStudy (calculusxy):

And how does this relate to the answer choices?

OpenStudy (calculusxy):

Can someone please help me right now? I need to finish my homework.

OpenStudy (calculusxy):

@TheSmartOne I can't really get what you are trying to say by saying that A and D look the same.

TheSmartOne (thesmartone):

Well A & D are exactly the same...

OpenStudy (calculusxy):

Sorry that was my typo. A has -4 as the sum

TheSmartOne (thesmartone):

ok lol

TheSmartOne (thesmartone):

well we coul first plug in (0,4) and see which ones it works for and elimnate 2 answer choices.

TheSmartOne (thesmartone):

could*

OpenStudy (calculusxy):

So you mean 0 for x and 4 for y right?

TheSmartOne (thesmartone):

Yes.

OpenStudy (calculusxy):

Okay so will it be D, then?

TheSmartOne (thesmartone):

Because if the y-intercept is 4 then that means x=0 and then y=4.

TheSmartOne (thesmartone):

When you plug that in you should get 2 choices... And then from there we can use the second peice of inf they gave us.

OpenStudy (calculusxy):

So is it not D because if x =0 and y=4, then 0+4=4?

OpenStudy (calculusxy):

@TheSmartOne

TheSmartOne (thesmartone):

It is D because 0+4=4 4=4 but there is 1 more answer choice

OpenStudy (calculusxy):

C?

TheSmartOne (thesmartone):

So it is Both C and D so far.. Now we need to use the second piece of informaton.

TheSmartOne (thesmartone):

Here is a graph of C and D. https://www.desmos.com/calculator/istu6qhknf Now which one slants downward @calculusxy @~

OpenStudy (calculusxy):

What do mean by the "second piece of information?"

TheSmartOne (thesmartone):

slants downward means a negative slope

OpenStudy (calculusxy):

Do you mean to say "left to right?"

TheSmartOne (thesmartone):

1st piece of information: slants downward 2nd piece of information: y-intercept is 4

OpenStudy (calculusxy):

The sum should be -4?

TheSmartOne (thesmartone):

their sum should be 4... remember 0+4=4 and 4-0=4

TheSmartOne (thesmartone):

slope is negative or slants downward means the line should be like this

OpenStudy (calculusxy):

so how do we narrow from C and D

OpenStudy (calculusxy):

i am not too good with slopes so i am having a problem...sorry if i am causing any disturbances :(

TheSmartOne (thesmartone):

well C has a positive slope and d has a negative slope...

TheSmartOne (thesmartone):

Here is the graph.. https://www.desmos.com/calculator/nhzcgep8lo

TheSmartOne (thesmartone):

SO it definitely ahs to be D.

TheSmartOne (thesmartone):

So*

OpenStudy (calculusxy):

Thank you so much!